# Generative Verifiers: Reward Modeling as Next-Token Prediction

### v1

@@ Abstract
Verifiers or reward models are often used to enhance the reasoning performance of large language models (LLMs). A common approach is the Best-of-N method, where N candidate solutions generated by the LLM are ranked by a verifier, and the best one is selected. While LLM-based verifiers are typically trained as discriminative classifiers to score solutions, they do not utilize the text generation capabilities of pretrained LLMs. To overcome this limitation, we instead propose training verifiers using the ubiquitous next-token prediction objective, jointly on verification and solution generation. Compared to standard verifiers, such generative verifiers (GenRM) can benefit from several advantages of LLMs: they integrate seamlessly with instruction tuning, enable chain-of-thought reasoning, and can utilize additional inference-time compute via majority voting for better verification. We demonstrate that when using Gemma-based verifiers on algorithmic and grade-school math reasoning tasks, GenRM outperforms discriminative verifiers and LLM-as-a-Judge, showing a $16-64$% improvement in the percentage of problems solved with Best-of-N. Furthermore, we show that GenRM scales favorably across dataset size, model capacity, and inference-time compute.
@@ 1 Introduction
Figure 1: Comparing GenRM with standard verification approaches on several reasoning tasks, in terms of Best-of-N performance. By recasting reward modeling as next-token prediction, GenRM outperforms other approaches to improve Best-of-N performance. The horizontal dotted line corresponds to direct generation performance (Best-of-1). GenRM leverages the generation capabilities of LLMs, enabling a finetuned verifier model to effectively utilize chain-of-thought and majority-voting at inference time to identify reasoning errors.
While large language models (LLMs) demonstrate remarkable capabilities, they often confidently make logical and factual mistakes (Zhang et al., 2023). These mistakes pose a significant challenge for reasoning problems, where a single mistake can invalidate the entire solution. A common strategy to address this issue is Best-of-N (Charniak and Johnson, 2005; Cobbe et al., 2021): the LLM generates N candidate solutions for a given problem, and a learned reward model, referred to as a “verifier”, ranks these solutions and picks the most suitable one. The effectiveness of this strategy hinges on how accurate the verifier is, making it crucial to identify better approaches for training verifiers.
On reasoning domains, LLM-based verifiers are typically trained as discriminative reward models (RMs) to assign numerical scores to candidate solutions, which is then used to classify them as correct or incorrect (Cobbe et al., 2021; Lightman et al., 2023; Wang et al., 2023). However, this scoring approach does not utilize the text-generation capabilities LLMs are fundamentally designed for. As a result, discriminative RMs miss out on the inherent strengths of generative LLMs, such as unified instruction tuning (Chung et al., 2022), chain-of-thought reasoning (Wei et al., 2022), and utilizing additional inference-time computation for better performance (Wang et al., 2022; Brown et al., 2024). While LLM-as-a-Judge (Zheng et al., 2024), which simply prompts off-the-shelf generative LLMs, also offer the above advantages, it often underperforms trained LLMs-based verifiers, especially on reasoning.
@@
Solution: Tim pays \$60 for cable. He gets Netflix for $10$ and the bundle of Hulu and Disney Plus costs \$10 * 80% = \$8. So he pays \$10 + \$8 = \$18 for Netflix and the bundle. Now he saves \$60 - \$18 = \$42. The answer is 42.
@@
GenRM-CoT: Let’s verify step by step.
$\cdots$
@@
Figure 2: Example using generative verifiers on GSM8K test. LLM-generated solutions often sound convincing even when they are wrong, making verification a challenging task. Here, the solution is incorrect because it has ignored the word ‘each’ in the problem. While the discriminative RM fails to recognize this subtle mistake in the solution, our GenRM-CoT verifier reliably detects the error. This is because GenRM-CoT was trained with next-token prediction on synthetic chain-of-thought rationales, enabling it to explicitly reason about the solution. The full verification output can be found in Table C.12.
In this work, we propose training verifiers with next-token prediction, which we call GenRM, to leverage the text generation capabilities of LLMs (Figure 2). Concretely, to produce a numerical score for a solution, the verifier now uses a prompt such as ‘Is the answer correct?’, and represents the score as the probability of a single text token (e.g., ‘Yes’ or ‘No’) under the context and the prompt. GenRM naturally supports CoT reasoning (Wei et al., 2022): it can be trained to reason explicitly by generating a verbalized rationale before predicting correctness using ‘Yes’ or ‘No’ token (Figure 3), assuming rationales are available during training. We can further boost verification accuracy of CoT verifiers using majority voting (Wang et al., 2022): sampling multiple CoT rationales (votes) and calculating the average score of the ‘Yes’ token across all samples, resulting in a favorable use of more test-time computation. Moreover, GenRM’s next-token prediction training enables unifying solution generation with verification, which has been difficult (Hosseini et al., 2024), potentially improving verification through positive knowledge transfer from solution generation.
GenRM outperforms discriminative RMs, LLM-as-a-Judge, and self-consistency on algorithmic string manipulation and math reasoning tasks (Figure 1), namely Last Letter Concat (Wei et al., 2022), Word Sorting from Big-Bench Hard (Suzgun et al., 2022), and GSM8K (Cobbe et al., 2021). Best-of-N performance with GenRM-CoT further improves when using majority-voting, nearly matching performance with oracle verifier on algorithmic reasoning tasks. On GSM8K (Figure 1, right), when using a Gemma-9B GenRM-CoT model to verify the outputs of Gemini 1.0 Pro, we observe a 20% improvement in terms of the number of problems solved ($73\%\rightarrow 92.8\%$), surpassing GPT-4 and Gemini 1.5 Pro. Moreover, we find that generative verifiers exhibit scales favorably as we increase dataset size as well as model capacity. Furthermore, GenRM-CoT also outperforms LLM-as-a-Judge as we scale inference-time compute by sampling multiple verification rationales for majority voting. Overall, these results suggest that generative verifiers hold significant potential for improving the reasoning capabilities of LLMs.
@@ 2 Preliminaries
$p_{\theta}({\mathbf{y}}\mid{\mathbf{x}})=\prod_{t=1}^{T}p_{\theta}(y_{t}\mid{\mathbf{x}},y_{<t}),$ (1)
with the convention $y_{<1}=\emptyset$ and ${\mathbf{y}}_{<t}=(y_{1},y_{2},\ldots,y_{t-1})$. For ease of notation, we define $p_{\theta}(y_{t}\mid{\mathbf{x}}):=p_{\theta}(y_{t}\mid{\mathbf{y}}_{<t},{\mathbf{x}})$. For a vocabulary size $M$, the probability of predicting the $t$-th token $y_{t}$, $p_{\theta}(y_{t}\mid{\mathbf{x}})$, is determined using a softmax with temperature $\gamma$ on logit scores $z$ of all the tokens: $p_{\theta}(y_{t}\mid{\mathbf{x}})=\frac{\exp(z_{t}/\gamma)}{\sum_{i=1}^{M}\exp(z_{i}/\gamma)},\quad\textrm{where}\ z_{t}=\mathrm{logit}_{\theta}(y_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t})$. Higher values of $\gamma$ introduce more randomness, while setting temperature $\tau=0$ makes the output deterministic, which corresponds to greedy decoding.
@@
${\mathcal{L}}_{\text{Discriminative-RM}}(\theta,{\mathcal{D}}_{RM})=$ $-\mathbb{E}_{({\mathbf{x}},{\mathbf{y}}^{+})\sim{\mathcal{D}}_{correct}}\left[\log r_{\theta}({\mathbf{x}},{\mathbf{y}}^{+})\right]-\mathbb{E}_{({\mathbf{x}},{\mathbf{y}}^{-})\sim{\mathcal{D}}_{incorrect}}\left[\log(1-r_{\theta}({\mathbf{x}},{\mathbf{y}}^{-}))\right],$ (3)
where $r_{\theta}({\mathbf{x}},{\mathbf{y}})=\text{sigmoid}(z_{cls}),\quad\text{and}\quad z_{cls}=\mathrm{logit}_{\theta}(cls\mid{\mathbf{y}},{\mathbf{x}})$
@@ 3 GenRM: Verification as Next-Token Prediction
Discriminative LLM-based verifiers (3) do not utilize the text generation capabilities of pretrained LLMs. To address this issue, we propose training verifiers or that can generate text, which we call GenRM, using standard next-token prediction (2). To do so, GenRM represents solution correctness using the LLM’s probability distribution over tokens, instead of predicting a separate numerical score. This keeps the generation abilities of GenRM intact as the verification decision is just another token, while also enabling several advantages that come for “free” with LLMs such as unified training for solution generation and verification, chain-of-thought reasoning, and inference-time computation.
@@ 3.1 Direct Verifier
In its simplest form, GenRM predicts whether a solution is correct using a single ‘Yes’ or ‘No’ token (Figure 3, top). This can be done by simply maximizing $\log p_{\theta}(\text{`Yes'}\mid({\mathbf{x}},{\mathbf{y}}^{+}))$ for correct solutions ${\mathbf{y}}^{+}$ and $\log p_{\theta}(\text{`No'}\mid({\mathbf{x}},{\mathbf{y}}^{-}))$ for incorrect solutions ${\mathbf{y}}^{-}$. To do so, we minimize the SFT loss in (2) on the dataset ${\mathcal{D}}_{\mathrm{Direct}}$ containing problem-solution pairs as input and a single ‘Yes‘ or ‘No’ token as target:
@@
This score takes into account the verifier’s confidence about its correctness prediction, which reduces the chance of being miscalibrated and wrong at test-time when using a binary ‘Yes’ or ‘No’ prediction.
@@ 3.2 Unifying Generation and Verification
GenRM seamlessly integrates reward modelling, which distinguishes between correct and incorrect solutions, with SFT for generating correct solutions. This can be done by simply changing the data mixture in the SFT loss (2) to include both verification and generation tasks. Given a verification dataset ${\mathcal{D}}_{\text{verify}}$, which can be ${\mathcal{D}}_{\text{Direct}}$ or ${\mathcal{D}}_{\text{CoT}}$ (discussed below) of problems-solution pairs with correctness tokens (optionally with CoT rationales), GenRM minimizes the loss:
@@
where $\lambda>0$ is a hyperparameter that controls the data mixture ratio between solution verification (${\mathcal{D}}_{\text{verify}}$) and generating correct solutions (${\mathcal{D}}_{\text{correct}}$). This unified training can improve verifier and generation performance via positive transfer between these two related tasks: how to generate a correct solution, and whether a solution is correct. By default, we train GenRM verifiers using the unified loss in (5).
@@ 3.3 Chain-of-Thought Verifiers (GenRM-CoT)
Since verification often involves complex reasoning, generative verifiers can naturally benefit from CoT reasoning (Wei et al., 2022). Specifically, we can generate intermediate reasoning steps or critique (CoT) before making a decision about the solution correctness, which may identify subtle reasoning errors missed by direct verifiers (Figure 3, bottom). To train CoT verifiers, we can minimize the SFT loss ${\mathcal{L}}_{{\text{GenRM}}}$ on the dataset ${\mathcal{D}}_{\text{CoT}}$ containing problem-solution pairs as inputs, and corresponding verification rationales ${\mathbf{v}}_{\textbf{CoT}}$ appended with a final question ${\mathbf{I}}$ and ‘Yes’ or ‘No’ token as targets:
@@
Notably, these rationales can either be human-generated or LLM-generated, both of which we explore in this work. During inference, we first generate a CoT rationale ${\mathbf{v}}_{\textbf{CoT}}$ from GenRM-CoT and then use the probability of ‘Yes’ for assigning the correctness score:
$r_{\text{CoT}}({\mathbf{x}},{\mathbf{y}})$ $=p_{\theta}(\text{Yes}\mid{\mathbf{x}},{\mathbf{y}},{\mathbf{I}}_{\textbf{CoT}},{\mathbf{v}}_{\textbf{CoT}},{\mathbf{I}}),\quad\text{where}\quad{\mathbf{v}}_{\textbf{CoT}}\sim p_{\theta}(\cdot\mid{\mathbf{x}},{\mathbf{y}},{\mathbf{I}}_{\textbf{CoT}}),$ (6)
Compared to (4) that directly uses the instruction ${\mathbf{I}}$ to produce a score, we can see that the above CoT reward additionally conditions on ${\mathbf{I}}_{\textbf{CoT}}$ and self-generated ${\mathbf{v}}_{\textbf{CoT}}$ before getting a score via instruction ${\mathbf{I}}$.
Inference-time Compute for CoT verifier When sampling verification CoTs, the generative verifier may use different reasoning paths and yield different correctness probabilities for the same problem-solution pair. As such, we would like to marginalize out the intermediate reasoning paths to select the most consistent correctness answer (Wang et al., 2022). To do so, we can use majority voting where we first generate $K$ verification CoT rationales, and average the CoT-verifier score for these rationales:
$r_{\text{MajVote@K}}({\mathbf{x}},{\mathbf{y}})=\dfrac{1}{K}\sum_{i=1}^{K}p_{\theta}\left(\text{Yes}\mid{\mathbf{x}},{\mathbf{y}},{\mathbf{I}}_{\textbf{CoT}},{\mathbf{v}}_{\textbf{CoT}}^{(i)},{\mathbf{I}}\right),\quad\text{where}\quad{\mathbf{v}}_{\textbf{CoT}}^{(i)}\sim p_{\theta}(\cdot\mid{\mathbf{x}},{\mathbf{y}},{\mathbf{I}}_{\textbf{CoT}})$ (7)
@@
Synthetic Verification Rationales for CoT Verifier. Verifying LLM solutions with human-generated rationales can become increasingly expensive and challenging as LLMs surpass human reasoning abilities. To address this challenge, we explore using synthetically-generated rationales on GSM8K. One naive approach is to simply use the ‘Let’s verify step by step’ prompt given a problem-solution pair, and filter the generated rationales based on whether they correctly verify the correctness of a solution. However, such rationales are often of poor quality due to 50% accuracy from random guessing.

### v2

@@ Abstract
Verifiers or reward models are often used to enhance the reasoning performance of large language models (LLMs). A common approach is the Best-of-N method, where N candidate solutions generated by the LLM are ranked by a verifier, and the best one is selected. While LLM-based verifiers are typically trained as discriminative classifiers to score solutions, they do not utilize the text generation capabilities of pretrained LLMs. To overcome this limitation, we instead propose training verifiers using the ubiquitous next-token prediction objective, jointly on verification and solution generation. Compared to standard verifiers, such generative verifiers (GenRM) can benefit from several advantages of LLMs: they integrate seamlessly with instruction tuning, enable chain-of-thought reasoning, and can utilize additional test-time compute via majority voting for better verification. We demonstrate that GenRM outperforms discriminative, DPO verifiers, and LLM-as-a-Judge, resulting in a $16-40$% improvement in the number of problems solved with Best-of-N on algorithmic and math reasoning tasks. Furthermore, we find that training GenRM with synthetic verification rationales is sufficient to pick out subtle errors on math problems. Finally, we demonstrate that generative verifiers scale favorably with model size and inference-time compute.
@@ 1 Introduction
[…]
While large language models (LLMs) demonstrate remarkable capabilities, they often confidently make logical and factual mistakes (Zhang et al., 2023). These mistakes pose a significant challenge for reasoning problems, where a single mistake can invalidate the solution. A common strategy to address this issue is Best-of-N (Charniak and Johnson, 2005; Cobbe et al., 2021): the LLM generates N candidate solutions for a given problem, and a learned reward model, referred to as a “verifier”, ranks these solutions and picks the most suitable one. The effectiveness of this strategy hinges on how accurate the verifier is, making it crucial to identify better approaches for training verifiers.
LLM-based verifiers for reasoning are typically trained as discriminative reward models (RMs) to assign numerical scores to candidate solutions, which is then used to classify them as correct or incorrect (Cobbe et al., 2021; Lightman et al., 2023; Wang et al., 2023). However, this scoring approach does not utilize the text-generation capabilities that LLMs are fundamentally designed for. As a result, discriminative RMs miss out on the inherent strengths of generative LLMs, such as unified instruction tuning (Chung et al., 2022), chain-of-thought (CoT) reasoning (Wei et al., 2022), and utilizing additional inference-time computation for better performance (Wang et al., 2022; Brown et al., 2024). While LLM-as-a-Judge (Zheng et al., 2024), which simply prompts off-the-shelf generative LLMs, also offers the above advantages, it typically underperforms trained LLMs-based verifiers on reasoning tasks, which we also observe in Figure 1.
@@
Solution: Tim pays \$60 for cable. He gets Netflix for $10$ and the bundle of Hulu and Disney Plus costs \$10 * 80% = \$8. So he pays \$10 + \$8 = \$18 for the bundle. Now he saves \$60 - \$18 = \$42. The answer is 42.
@@
GenRM-CoT: Let’s verify step by step. $\cdots$
@@
Figure 2: Example using generative CoT verifier on GSM8K test. LLM-generated solutions often sound convincing even when they are wrong, making verification a challenging task. Here, the solution is incorrect because it has ignored the word ‘each’ in the problem. While the discriminative RM fails to recognize this subtle mistake in the solution, our GenRM-CoT verifier reliably detects the error. This is because GenRM-CoT was trained with next-token prediction on synthetic chain-of-thought rationales, enabling it to explicitly reason about the solution. The full verification output can be found in Table E.12.
In this work, we propose training verifiers with next-token prediction, which we call GenRM, to leverage the text generation capabilities of LLMs (Figure 2). Concretely, to produce a numerical score for a solution, the verifier now uses a prompt such as ‘Is the answer correct?’, and represents the score as the probability of a single text token (e.g., ‘Yes’ or ‘No’). GenRM naturally supports CoT reasoning (Wei et al., 2022): it can be trained to reason explicitly by generating a verbalized rationale before predicting correctness using ‘Yes’ or ‘No’ token (Figure 3), assuming rationales are available during training. We can further boost verification accuracy of CoT verifiers using majority-voting (Wang et al., 2022): sampling multiple CoT rationales and calculating the average score of the ‘Yes’ token across all rationales, enabling the use of inference-time compute for verification. Moreover, GenRM’s next-token prediction training enables unifying solution generation with verification, which has been difficult with DPO verifiers (Rafailov et al., 2024; Hosseini et al., 2024), potentially improving verification through positive transfer from solution.
Our results show that GenRM outperforms discriminative RMs, LLM-as-a-Judge, and self-consistency on algorithmic string manipulation and math reasoning tasks (Figure 1). Best-of-N performance further improves with GenRM-CoT that uses majority-voting, nearly matching performance with oracle verifier on algorithmic tasks. On GSM8K, when using a Gemma2-9B GenRM-CoT verifier on solutions from Gemini 1.0 Pro, we observe an improvement from $73\%\rightarrow 93.4\%$ in terms of the number of problems solved, surpassing GPT-4 and Gemini 1.5 Pro. Furthermore, GenRM-CoT trained on grade-school math problems exhibit easy-to-hard generalization, solving 17% more high-school competition problems in MATH500 (Lightman et al., 2023) with Best-of-32. Moreover, we find that generative verifiers scale more favorably than discriminative verifiers as we increase model capacity, and outperform LLM-as-a-Judge as we scale inference-time compute with majority voting. Overall, generative verifiers hold significant potential for improving the reasoning capabilities of LLMs.
@@ 2 Preliminaries
$p_{\theta}({\mathbf{y}}\mid{\mathbf{x}})=\prod_{t=1}^{T}p_{\theta}(y_{t}\mid{\mathbf{x}},y_{<t})$ (1)
with the convention $y_{<1}=\emptyset$ and ${\mathbf{y}}_{<t}=(y_{1},y_{2},\ldots,y_{t-1})$. For ease of notation, we define $p_{\theta}(y_{t}\mid{\mathbf{x}}):=p_{\theta}(y_{t}\mid{\mathbf{y}}_{<t},{\mathbf{x}})$. For a vocabulary size $M$, the probability of predicting the $t$-th token $y_{t}$, $p_{\theta}(y_{t}\mid{\mathbf{x}})$, is determined using a softmax with temperature $\gamma$ on logit scores $z$ of all the tokens: $p_{\theta}(y_{t}\mid{\mathbf{x}})=\frac{\exp(z_{t}/\gamma)}{\sum_{i=1}^{M}\exp(z_{i}/\gamma)},\quad\textrm{where}\ z_{t}=\mathrm{logit}_{\theta}(y_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t})$. Higher values of temperature $\gamma$ introduce more randomness, while setting temperature $\tau=0$ makes the output deterministic, which corresponds to greedy decoding.
@@
${\mathcal{L}}(\theta,{\mathcal{D}}_{RM})=$ $-\mathbb{E}_{({\mathbf{x}},{\mathbf{y}}^{+})\sim{\mathcal{D}}_{\text{correct}}}\left[\log r_{\theta}({\mathbf{x}},{\mathbf{y}}^{+})\right]-\mathbb{E}_{({\mathbf{x}},{\mathbf{y}}^{-})\sim{\mathcal{D}}_{\text{incorrect}}}\left[\log(1-r_{\theta}({\mathbf{x}},{\mathbf{y}}^{-}))\right],$
where $r_{\theta}({\mathbf{x}},{\mathbf{y}})=\text{sigmoid}(z_{cls}),\quad\text{and}\quad z_{cls}=\mathrm{logit}_{\theta}(cls\mid{\mathbf{y}},{\mathbf{x}})$ (3)
@@ 3 GenRM: Verification as Next-Token Prediction
Discriminative LLM-based verifiers (3) do not utilize the text generation capabilities of pretrained LLMs. To address this issue, we propose training generative verifiers, which we call GenRM, using standard next-token prediction (2). To do so, GenRM represents solution correctness using the LLM’s probability distribution over tokens, instead of predicting a separate numerical score. This keeps the generation abilities of GenRM intact as the verification decision is just another token, while also enabling several advantages that come for “free” with LLMs, such as unified training for solution generation and verification, chain-of-thought reasoning, and inference-time computation.
@@ 3.1 Direct Verifier
In its simplest form, GenRM predicts whether a solution is correct using a single ‘Yes’ or ‘No’ token (Figure 3, top). This can be done by maximizing $\log p_{\theta}(\text{`Yes'}\mid({\mathbf{x}},{\mathbf{y}}^{+}))$ for correct solutions ${\mathbf{y}}^{+}$ and $\log p_{\theta}(\text{`No'}\mid({\mathbf{x}},{\mathbf{y}}^{-}))$ for incorrect solutions ${\mathbf{y}}^{-}$. To do so, we minimize the SFT loss in (2) on the dataset ${\mathcal{D}}_{\mathrm{Direct}}$ containing problem-solution pairs and a ‘Yes‘ or ‘No’ verification token:
@@
This score takes into account the verifier’s confidence about its correctness prediction, which reduces the chance of being wrong at test-time when using a binary ‘Yes’ or ‘No’ prediction.
@@ 3.2 Unifying Generation and Verification
GenRM seamlessly integrates reward modeling, which distinguishes between correct and incorrect solutions, with SFT for generating correct solutions. This can be done by simply changing the data mixture in the SFT loss (2) to include both verification and generation tasks. Given a verification dataset ${\mathcal{D}}_{\text{verify}}$, which can be ${\mathcal{D}}_{\text{Direct}}$ or ${\mathcal{D}}_{\text{CoT}}$ (discussed below) of problems-solution pairs with correctness tokens (optionally with CoT rationales), GenRM minimizes the loss:
@@
where $\lambda>0$ is a hyperparameter that controls the mixture ratio between verification (${\mathcal{D}}_{\text{verify}}$) and generating correct solutions (${\mathcal{D}}_{\text{correct}}$). This unified training can improve verifier and generation performance via positive transfer between these two related tasks: how to generate a correct solution, and whether a solution is correct. By default, we train GenRM verifiers using the unified loss in (5).
@@ 3.3 Chain-of-Thought Verifiers (GenRM-CoT)
Since verification often involves nuanced reasoning, generative verifiers can naturally benefit from CoT (Wei et al., 2022). Specifically, we can generate intermediate reasoning steps or critique (CoT) before making a decision about the solution correctness, which may identify subtle reasoning errors missed by direct verifiers (Figure 3, bottom). To train CoT verifiers, we can minimize the SFT loss ${\mathcal{L}}_{{\text{GenRM}}}$ on the dataset ${\mathcal{D}}_{\text{CoT}}$ containing problem-solution pairs as inputs, and corresponding verification rationales ${\mathbf{v}}_{\textbf{CoT}}$ appended with a final question ${\mathbf{I}}$ and ‘Yes’ or ‘No’ token as targets:
@@
where ${\mathbf{I}}_{\textbf{CoT}}=$‘Let’s verify step by step.’. Notably, these rationales can either be human or LLM-generated, both of which we explore in this work. During inference, we first generate a CoT rationale ${\mathbf{v}}_{\textbf{CoT}}$ from GenRM-CoT and then use the probability of ‘Yes’ for assigning the correctness score:
$r_{\text{CoT}}({\mathbf{x}},{\mathbf{y}})$ $=p_{\theta}(\text{Yes}\mid{\mathbf{x}},{\mathbf{y}},{\mathbf{I}}_{\textbf{CoT}},{\mathbf{v}}_{\textbf{CoT}},{\mathbf{I}}),\quad\text{where}\ \ {\mathbf{v}}_{\textbf{CoT}}\sim p_{\theta}(\cdot\mid{\mathbf{x}},{\mathbf{y}},{\mathbf{I}}_{\textbf{CoT}}),$ (6)
Compared to (4) that only uses the instruction ${\mathbf{I}}$ to produce a score, the above CoT reward additionally conditions on ${\mathbf{I}}_{\textbf{CoT}}$ and self-generated ${\mathbf{v}}_{\textbf{CoT}}$ before getting a score via instruction ${\mathbf{I}}$.
Inference-time compute for CoT verifier. When sampling verification CoTs, the generative verifier can use different reasoning paths and yield different correctness probabilities for the same problem-solution pair. As such, we would like to marginalize out these reasoning paths to select the most consistent correctness answer (Wang et al., 2022). To do so, we use majority voting where we first generate $K$ verification CoT rationales, and average the CoT-verifier score for these rationales:
$r_{\text{MajV@K}}({\mathbf{x}},{\mathbf{y}})=\dfrac{1}{K}\sum_{i=1}^{K}p_{\theta}\left(\text{Yes}\mid{\mathbf{x}},{\mathbf{y}},{\mathbf{I}}_{\textbf{CoT}},{\mathbf{v}}_{\textbf{CoT}}^{(i)},{\mathbf{I}}\right),\quad\text{where}\ \ {\mathbf{v}}_{\textbf{CoT}}^{(i)}\sim p_{\theta}(\cdot\mid{\mathbf{x}},{\mathbf{y}},{\mathbf{I}}_{\textbf{CoT}})$ (7)
@@
Synthetic verification CoT rationales for training. Verifying LLM solutions with human-generated rationales can become increasingly expensive and challenging as LLMs surpass human reasoning abilities. To address this challenge, we explore using synthetically-generated rationales on GSM8K. One naive approach is to simply use the ‘Let’s verify step by step’ prompt given a problem-solution pair, and keep the generated rationales only when they accurately verify the correctness of a solution (Singh et al., 2023; Zelikman et al., 2022). However, such rationales (after filtering based on final yes/no responses) are still often of poor quality, due to 50% accuracy from random guessing.